OpenStudy (anonymous):

What is the common difference of a 43–term arithmetic sequence where the first term is –13 and the sum is 9,374?

OpenStudy (anonymous):

S43 = 9374 = (2(13) + (42)d)(43/2)

OpenStudy (anonymous):

Each term adds that common difference once, if there are 43 terms, then that difference has been added 42 times. -13 + 42D = 9374 where D is the common difference

OpenStudy (anonymous):

No Smooth. The total is 9374, not the last term

OpenStudy (anonymous):

Ah. Indeed.

OpenStudy (anonymous):

Then in that case, ((-13+42D) + -13)*43/2 = 9374

OpenStudy (anonymous):

so D = 11

OpenStudy (apoorvk):

Sum 'S' of an AP of 'n' terms with the first term 'a' and common difference 'd' is: \[\large S = \frac n 2 [2a + (n-1)d]\] So, plug in the values of 'S', 'n' and 'a', calculate 'd' and enjoy!

OpenStudy (anonymous):

-13. ((-13+42D) + -13)*43/2 = 9374 gives (42D -26)*43/2 = 9374 42d-26 = 9374*(2/43) 42D -26= 436 42D = 462 D = 462/42 = 11

OpenStudy (anonymous):

how do i plug this one in "What is the sum of a 14–term arithmetic sequence where the last term is 30 and the common difference is –5?"

OpenStudy (anonymous):

The formula for the sum is: [(First term) + (Last term) ]* (number of terms divided by 2) You're given the last term. You can calculate the first term by (First term) -(D*n-1) where n is the number of terms. Plug in and solve.

OpenStudy (anonymous):

812.5?

OpenStudy (apoorvk):

in the equation i wrote above, the part withing the brackets can be written as ...[a + a+(n-1)d] here, if you notice, a+(n-1)d is actually the last term 'l'. so you can rewrite the whole formula as: \[\large S= \frac n 2 (a+l)\]

OpenStudy (apoorvk):

okay when its given that the last term is given, like for example a, b, c, d, ..... , y,z, 25 reverse the series and write 25, y, z, ...., d, c, b, a So now the first term becomes 25. And ofcourse the common difference which was 'd' earlier, becomes '-d' (since the AP has been reversed -if it was an increasing one originally it is now decreasing, and vice-versa). And the no. of terms 'n' remains the same. Now do the calculation for this. and then adjust for the original AP accordingly.